OpenStudy (anonymous):

If f(x) = x - 2/ 2, then what is f^–1 (x) ?

OpenStudy (anonymous):

Replace x and y to each other then solve for y.

OpenStudy (anonymous):

f(x) = y = x - 2 / 2, so f^(-1)(x) = x = y - 2 / 2 Now solve for y.

OpenStudy (anonymous):

x = (y - 2) / 2 2x = y - 2 2x + 2 = y right?

OpenStudy (anonymous):